• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:03:15 on March 17, 2006

    Entry number 166969

    keyword=Beam Time Accounting
    Hall A

    03/16/06 1600-2400
    Experiment #: E02-013
    Shift Leader: Vina Punjabi
    Shift Workers: Klaus Grimm

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.000.000.000.000.000.000.000.000.00
    Beam Available
    but Not in Use
    0.570.870.940.980.961.000.971.007.29
    Beam Not Available
    or Unacceptable
    0.430.130.060.020.040.000.030.000.71
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    0.430.000.000.000.000.000.000.000.43
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.571.001.001.001.001.001.001.007.57
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.571.001.001.001.001.001.001.007.57
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: During controlled access to fix magnet problem for bigbite, Doug discovered problem of leak in Be window. Problem will be fixed in the morning. So no beam in swing shift.

    Data Entry by: Vina Pnjabi, Klaus Grimm